a goldsmith combined an alloy that cost 4.10 per ounce with an alloy that cost 1.80 per ounce. how many ounces of each were used to make of 200 oz costing 3.41 per ounce.
Answer by stanbon(75887)

a goldsmith combined an alloy that cost 4.10 per ounce with an alloy that cost 1.80 per ounce. how many ounces of each were used to make of 200 oz costing 3.41 per ounce.
------
Value Equation:
4.1x + 1.8(200-x) = 3.41*200
------
Multiply thru by 100 to get:
410x + 180*200 - 180x = 341*200
------
230x = 161*200
------
x = 140 oz (amt. of $4.10 alloy needed)
---
200-x = 60 oz (amt. of $1.80 alloy needed)
==============================================
Cheers,
Stan H.
===================